If f(x)=x+4 and g(x)=2x^2-x-1, find the composition (fog)(2)
Please help me I get so far and then I can not get the rest. Thank you

Answer by scott8148(6628)

(fog)(x) is f(g(x))

f(2x^2-x-1)=(2x^2-x-1)+4

(fog)(2)=(2(2^2)-2-1)+4 __ (fog)(2)=(5)+4